Check the extension of d given by d(x,y) = d(x,a) + 1 + d(b,y) where x belongs to X and y belongs to Y/ X
1
Expert's answer
2021-09-21T03:04:13-0400

Given that

"d(x,y) = d(x,a) + 1 + d(b,y) \\
x \in X, y \in \frac{Y}{X}\\
So, y = x_0+y_0 \space s.t \space x_0 \in x, y_0 \in y\\
\implies d(x,y) - d(x,a) = 1 + d(b,y) \\
d(y,a) ≤ d(x,y) -d(x,a) = 1 + d(b,y) \\
d(y,a) -d(b,y) ≤1 \\
d(y,a) -d(y,b) ≤1 \\
d(a,b) ≤ d(y,a) -d(y,b)≤1\\
d(a,b)≤1 \space for \space all \space a,b"